OpenStudy (gavin39):

need quick question answered If f(x) = log3 (x + 1), what is f−1(2)? 1 8 10 27

OpenStudy (gavin39):

I give medals and fan

hero (hero):

\(f(x) = log_3(x + 1)\) Let y = f(x) \(y = log_3(x + 1)\) Swap x and y: \(x = log_3(y + 1)\) Apply the "relationship" rule for logs and exponents: \(3^x = y + 1\) Subtract 1 from both sides: \(3^x - 1 = y\) Replace y with \(f^{-1}(x)\) \(3^x - 1 = f^{-1}(x)\) Rewrite as \(f^{-1}(x) = 3^x - 1\) Can you now try evaluating \(f^{-1}(2)\)?

OpenStudy (gavin39):

in all honesty I cant because I know nothing about this stuff. Confuses me bady

hero (hero):

I see...

OpenStudy (gavin39):

O wait I got it. Its 8

hero (hero):

Yes, but now, try to understand the entire process.

OpenStudy (gavin39):

yeah I need to study hard on it

hero (hero):

There are only a few things that might trip you up: 1. Replacing f(x) with y 2. Swapping x and y 3. Applying the relationship rule for logs and exponents. 4. Isolating y 5. Replacing y with f^{-1}(x) Anytime you have to find the inverse of a function, you'll have to do such things.

hero (hero):

Step 3 will be applied at certain points when dealing with finding the inverse of logs or exponents. All the other steps are general when finding the inverse of any function with x and y as variables.